A basketball spinning on someone’s finger undergoes an angular acceleration of - 0.15 rad/s and is traveling at 27 rad/s initially. How many revolutions will it make before coming to a stop?


1
Expert's answer
2021-01-12T12:13:21-0500

Let's first find time that the basketball takes to stop:


"\omega=\omega_0+\alpha t,""t=-\dfrac{\omega_0}{\alpha}=-\dfrac{27\ \dfrac{rad}{s}}{-0.15\ \dfrac{rad}{s^2}}=180\ s."


Then, we can find how many radians the basketball turns through:


"\theta=\omega_0 t+\dfrac{1}{2}\alpha t^2,""\theta=27\ \dfrac{rad}{s}\cdot 180\ s+\dfrac{1}{2}\cdot(-0.15\ \dfrac{rad}{s^2})\cdot(180\ s)^2=2430\ rad."

Finally, we can find the number of revolutions that the basketball makes before coming to a stop:


"N_{rev}=2430\ rad\cdot(\dfrac{1\ rev}{2\pi\ rad})=387\ rev."

Answer:

"N_{rev}=387\ rev."